OpenStudy (anonymous):

I need an URGENT answer PLEASE! Rukia mixed 2 holds 5/18 litres of blue paint with 1 holds 5/15 litres of white paint to make a sky blue paint. How many litres of sky blue paint did she make?

OpenStudy (anonymous):

blue paint (litres) = 2*5/18 = 5/9 white paint (litres) = 5/15 = 1/3 Together: 5/9+1/3=5/9+3/9=8/9 litres